# A universal neutral-atom quantum computer with individual optical addressing and non-destructive readout

### v1

@@ 1 Introduction
Quantum computers promise to reduce computational energy consumption and solve classically expensive or intractable problems in a range of fields including material science, machine learning, security, chemistry, energy, finance, medicine, and pharmaceutical development [1, 2, 3, 4, 11, 12, 13, 14, 15]. Recent analyses suggest that realization of practical utility will require fault-tolerant operation [5]. This in turn requires quantum error correction (QEC), which comes with large overhead in the number of qubits and requires high one- and two-qubit quantum gate fidelities. The field’s progress along this path has entered a phase of experimental demonstrations [16, 17, 7]. Among the diverse modalities proposed for large-scale fault-tolerant quantum computation [18], the neutral-atom approach is one of the most promising candidates and has become a leader in the combination of qubit count, universal gate-based circuit operation, connectivity, and quantum error detection [7, 19, 20]. The rapid progress of neutral atoms for quantum computing can be attributed to the combination of naturally identical qubits; long coherence times due to decoupling from the environment [21]; switchable qubit interaction strengths across 12 orders of magnitude [22]; sub-microsecond, high-fidelity, scalable gates [6, 23]; and the ability to trap and manipulate large, dense, multi-dimensional qubit arrays [24, 19, 25], as was recently highlighted with a demonstration of error-detected logical qubits [7].
@@ 3 Quantum processor implementation and characterization
Figure 2: Single-qubit gates. a, Rabi oscillations driven by a global microwave field. b, Ramsey oscillations from local qubit $R_{Z}(\phi)$ control. c, d, Randomized benchmarking results for global $R(\theta,\phi)$ and local $R_{Z}(\phi)$ gates, respectively. e, Randomized benchmarking results, plotted as cumulative distribution function (CDF) of 24-qubit (or 1-qubit) results. For 24-qubit results, gates are applied to each of the 24 qubits in each circuit.
@@ 4 Non-destructive State Selective Readout
Figure 4: Non-destructive state-selective readout. NDSSR signal distribution from 51862 images collected during 43 iterations of SPAM characterization. The turquoise-colored distribution describes the signals where atoms are expected to be in a bright state ($|F=4\rangle$) and the purple-colored distribution describes the signals for a dark state ($|F=3\rangle$). The samples for each distribution come from the deterministically loaded diagnostic site pair used for CZ gate fidelity benchmarking. Photo-electron counts for each site were calculated from the raw EMCCD camera analog-digital units using nominal EM gain and pre-amplifier gain. Each quadrant of the plane defined by NDSSR and post-NDSSR photo-electron counts contains two text labels showing total counts from $|F=4\rangle$ preparation and measurement experiment (turquoise) and counts from $|F=3\rangle$ preparation and measurement experiment (purple). a, Sampled probability distribution of photo-electron counts for bright and dark states. The dashed black line denotes the discrimination threshold that yields discrimination fidelity of 99.6(2)%. The shaded areas under the distributions on both sides of the discrimination threshold denote signals associated with state detection error. Depumping and atom loss during the NDSSR process and microwave pulse infidelity during state transfer from the stretched state to the bright qubit state give rise to bright state detection error of 3.1(7)%, which is reduced to 2.6(7)when atom loss is post-selected out. The dark state detection error of 1.6(5)is believed to be a combination of (partial) state preparation error and $F=3\rightarrow F=4$ repumping during imaging from residual leakage of repump laser intensity. The middle inset shows a single NDSSR measurement of a four-qubit GHZ state realized in the 24-qubit array. b, NDSSR and post-NDSSR occupancy readout photo-electron counts. The bright state atom loss rate is 1.0(4)and the dark state atom loss rate is 0.8(3)%. The atom loss rates and state detection errors provided are the means and standard deviation of their measured values for the 43 experimental iterations that comprise the dataset.

### v2

@@ 1 Introduction
Quantum computers promise to reduce computational energy consumption and solve classically expensive or intractable problems in a range of fields including material science, machine learning, security, chemistry, energy, finance, medicine, and pharmaceutical development [1, 2, 3, 4, 11, 12, 13, 14, 15]. Recent analyses suggest that realization of practical utility will require fault-tolerant operation [5]. This in turn requires quantum error correction (QEC), which comes with large overhead in the number of qubits and requires high one- and two-qubit quantum gate fidelities. The field’s progress along this path has entered a phase of experimental demonstrations [16, 17, 7]. Among the diverse modalities proposed for large-scale fault-tolerant quantum computation [18], the neutral-atom approach is one of the most promising candidates and has become a leader in the combination of qubit count, universal gate-based circuit operation, connectivity, and quantum error correction [7, 19, 20]. The rapid progress of neutral atoms for quantum computing can be attributed to the combination of naturally identical qubits; long coherence times due to decoupling from the environment [21]; switchable qubit interaction strengths across 12 orders of magnitude [22]; sub-microsecond, high-fidelity, scalable gates [6, 23]; and the ability to trap and manipulate large, dense, multi-dimensional qubit arrays [24, 19, 25], as was recently highlighted with a demonstration of error-corrected logical qubits [7].
@@ 3 Quantum processor implementation and characterization
Figure 2: Single-qubit gates. a, Rabi oscillations driven by a global microwave field. b, Ramsey oscillations from local qubit $R_{Z}(\phi)$ control. c, d, Randomized benchmarking results for global $R(\theta,\phi)$ and local $R_{Z}(\phi)$ gates, respectively. Each point is an average over multiple random circuits (10 per depth for $R(\theta,\phi)$ and 5 per depth for $R_{Z}(\phi)$ ), with error bar given by standard error of the mean. e, Randomized benchmarking results, plotted as cumulative distribution function (CDF) of 24-qubit (or 1-qubit) results. For 24-qubit results, gates are applied to each of the 24 qubits in each circuit. For the single qubit results, the reported uncertainty is the benchmarking fit uncertainty. For the 24-qubit results, the reported uncertainty is the ensemble standard deviation.
@@ 4 Non-destructive State Selective Readout
Figure 4: Non-destructive state-selective readout. NDSSR signal distribution from 51862 images collected during 43 iterations of SPAM characterization. The turquoise-colored distribution describes the signals where atoms are expected to be in a bright state ($|F=4\rangle$) and the purple-colored distribution describes the signals for a dark state ($|F=3\rangle$). The samples for each distribution come from the deterministically loaded diagnostic site pair used for CZ gate fidelity benchmarking. Photo-electron counts for each site were calculated from the raw EMCCD camera analog-digital units using nominal EM gain and pre-amplifier gain. Each quadrant of the plane defined by NDSSR and post-NDSSR photo-electron counts contains two text labels showing total counts from $|F=4\rangle$ preparation and measurement experiment (turquoise) and counts from $|F=3\rangle$ preparation and measurement experiment (purple). a, Sampled probability distribution of photo-electron counts for bright and dark states. The dashed black line denotes the discrimination threshold that yields discrimination fidelity of 99.6(2)%. The shaded areas under the distributions on both sides of the discrimination threshold denote signals associated with state detection error. Depumping and atom loss during the NDSSR process and microwave pulse infidelity during state transfer from the stretched state to the bright qubit state give rise to bright state detection error of 3.1(7)%, which is reduced to 2.6(7)when atom loss is post-selected out. The dark state detection error of 1.6(5)is believed to be a combination of (partial) state preparation error and $F=3\rightarrow F=4$ repumping during imaging from residual leakage of repump laser intensity. The middle inset shows a single NDSSR measurement of a four-qubit GHZ state realized in the 24-qubit array. b, NDSSR and post-NDSSR occupancy readout photo-electron counts. The bright state atom loss rate is 1.0(4)and the dark state atom loss rate is 0.8(3)%. The atom loss rates and state detection errors and their uncertainties provided are the means and standard deviation of their measured values for the 43 experimental iterations that comprise the dataset.